### v1

Abstract
The rising popularity of multimodal large language models (MLLMs) has sparked a significant increase in research dedicated to evaluating these models. However, current evaluation studies predominantly concentrate on the ability of models to comprehend and reason within a unimodal (vision-only) context, overlooking critical performance evaluations in complex multimodal reasoning tasks that integrate both visual and text contexts. Furthermore, tasks that demand reasoning across multiple modalities pose greater challenges and require a deep understanding of multimodal contexts. In this paper, we introduce a comprehensive assessment framework named MM-InstructEval, which integrates a diverse array of metrics to provide an extensive evaluation of the performance of various models and instructions across a broad range of multimodal reasoning tasks with vision-text contexts. MM-InstructEval enhances the research on the performance of MLLMs in complex multimodal reasoning tasks, facilitating a more thorough and holistic zero-shot evaluation of MLLMs. We firstly utilize the “Best Performance” metric to determine the upper performance limit of each model across various datasets. The “Mean Relative Gain” metric provides an analysis of the overall performance across different models and instructions, while the “Stability” metric evaluates their sensitivity to variations. Historically, the research has focused on evaluating models independently or solely assessing instructions, overlooking the interplay between models and instructions. To address this gap, we introduce the “Adaptability” metric, designed to quantify the degree of adaptability between models and instructions. Evaluations are conducted on 31 models (23 MLLMs) across 16 multimodal datasets, covering 6 tasks, with 10 distinct instructions. The extensive analysis enables us to derive novel insights.
Keywords:
Multimodal Large Language Models Multimodal Reasoning Tasks with Vision-Text Contexts Various Instructions Zero-shot Evaluation Comprehensive Metrics
Figure 1: Required capabilities for diverse datasets. Different from the traditional capabilities indicated above the dotted line for multimodal reasoning tasks with vision-only contexts—including COCO (Lin et al., 2014), VQA v2 (Lin et al., 2014), Text VQA (Singh et al., 2019), OK-VQA (Marino et al., 2019), and MM-VET (Yu et al., 2023b)—our own multimodal reasoning tasks that incorporate vision-text contexts, particularly those below the dotted line, require not only capabilities relevant to traditional tasks but also a profound interaction and understanding of complex vision-text contexts. In these tasks, ‘T’ represents the text context, ‘Q’ denotes the question prompting the models for answers, and ‘GT’ stands for the ground truth label. ‘HE’ and ‘TE’ correspond to the head entity and tail entity, respectively.
1 Introduction
Benchmark Purpose Text Context Image Question Answer Type Evaluator
MME (Fu et al., 2023) Traditional Capabilities Yes/No Metrics
MMBench (Liu et al., 2023c) Traditional Capabilities Multi-choice GPT
SEED-Bench (Li et al., 2023a) Traditional Capabilities Multi-choice Metrics
LVLM-eHub (Xu et al., 2023) Traditional Capabilities Multi-choice/Open-ended Metrics
MM-Vet (Yu et al., 2023b) Traditional Capabilities Multi-choice/Open-ended Metrics/GPT-4
MLLM-Bench (Ge et al., 2023) Traditional Capabilities Open-ended GPT-4V/LLaVA
MMMU (Yue et al., 2023) Multimodal Reasoning Multi-choice/Open-ended Metrics
HallusionBench (Guan et al., 2023) Visual Hallucination Yes/No Metrics
Bingo (Cui et al., 2023) Visual Hallucination Open-ended Human
MM-InstructEval Multimodal Reasoning/Instructions Multi-choice Metrics
Table 1: Comparison of our proposed MM-InstructEval with recent vision-language benchmarks. Traditional capabilities, as mentioned in previous studies (Yu et al., 2023b; Li et al., 2023a), include Language Generation, Recognition, Knowledge Reasoning, Spatial Reasoning, OCR, and more. Multimodal reasoning tasks with multimodal contexts more concentrate on Multimodal Alignment, Multimodal Fusion, Multimodal Content Understanding, and other abilities. MM-InstructEval also conducted a study about the adaptability between diverse models and instructions.
Multimodal Large Language Models (MLLMs), harnessing the formidable capabilities of Large Language Models (LLMs), demonstrate outstanding performance across a spectrum of multimodal tasks (Alayrac et al., 2022; Gao et al., 2023; Li et al., 2023b; Dai et al., 2023). Recent research developments, as summarized in Table 1, include but are not limited to MME (Fu et al., 2023), MMBench (Liu et al., 2023c), SEED-Bench (Li et al., 2023a), LVLM-eHub (Xu et al., 2023), MM-Vet (Yu et al., 2023b), MMMU (Yue et al., 2023), and MLLM-Bench (Ge et al., 2023). These studies primarily focus on evaluating the traditional vision-language multimodal capabilities of MLLMs in tasks predominantly driven by visual context (vision context + text question), such as Visual Question Answering (VQA) and Video Question Answering (VideoQA). In these tasks, the answers to the questions are often derived solely based on the image context, as illustrated in the section above the dotted line in Figure 1. These capabilities include Recognition, Optical Character Recognition (OCR), Spatial Reasoning, Knowledge Integration, and Code Reasoning (Yu et al., 2023b). However, the performance of Multimodal Large Language Models (MLLMs) in multimodal reasoning tasks that combine vision-text contexts (vision-text contexts + text question) remains less explored. Such tasks include Visual Question Answering with Multimodal Contexts (VQAMC) (Lu et al., 2022a; Yue et al., 2023; Ghosal et al., 2024), Multimodal Sentiment Analysis (MSA) (Niu et al., 2016; Yang et al., 2021a; Zadeh et al., 2016, 2018), Multimodal Aspect-Based Sentiment Analysis (MABSA) (Zhang et al., 2018; Lu et al., 2018; Zhou et al., 2021), Multimodal Hateful Memes Detection (MHMD) (Mathias et al., 2021), Multimodal Sarcasm Detection (MSD) (Cai et al., 2019), and Multimodal Relation Extraction (MRE) (Mathias et al., 2021). In the field of Natural Language Processing (NLP), most research (Zhang et al., 2023b; Wang et al., 2023; Dalvi et al., 2023) primarily evaluates pure Large Language Models (LLMs) like ChatGPT (OpenAI, 2023a), Flan-T5 (Chung et al., 2022), focusing specifically on text classification tasks such as text sentiment analysis and relation classification. It leaves a gap in understanding the performance of various MLLMs in multimodal reasoning tasks that rely on both text and image contexts. For example, in the Visual Question Answering with Multimodal Contexts (VQAMC) task depicted in Figure 1, it is clear that the answer to the question depends on integrating the image-text contexts. Simply using the image and the question without incorporating the textual context fails to produce the correct answer.
To address the aforementioned gap, we conduct the comprehensive evaluation involving 31 publicly available models, including 23 Multimodal Large Language Models (MLLMs), across a diverse set of 16 datasets covering 6 distinct tasks, as detailed in Tables 2 and 3. Our primary objective is to assess the performance of various MLLMs for multimodal reasoning tasks that integrate vision-text contexts. Additionally, we aim to establish benchmarks for a range of MLLMs across diverse multimodal reasoning tasks. These tasks demand not only traditional vision-language multimodal capabilities from the models but also require a deep understanding of multimodal contexts for effective reasoning to arrive at the final answer. It is illustrated in the section below the dotted line in Figure 1. Multimodal reasoning tasks with vision-text contexts often necessitate various cognitive processes, including but not limited to multimodal alignment, multimodal interaction, and multimodal fusion, as emphasized in studies (Han et al., 2021; Hazarika et al., 2020; Wu and Li, 2023), alongside content understanding (Wu et al., 2024). Defining the boundaries of such multimodal reasoning with vision-text contexts can be challenging, as the multimodal contexts may encompass diverse elements such as visual-language alignment through OCR and spatial reasoning, among others. Furthermore, in tasks like visual question answering with multimodal contexts, especially when accompanied by prompts or instructions, these reasoning abilities are often triggered, as vividly demonstrated in datasets like AlgoPuzzleVQA (Ghosal et al., 2024).
Figure 2: Overview of our MM-InstructEval framework, which conducts evaluations of popular MLLMs across various multimodal reasoning tasks with multimodal contexts utilizing comprehensive metrics. As illustrated in the dashed box, we select the ‘MVSA-Single’ dataset from the ‘MSC’ task to utilize ‘Instruction # 2’ for evaluating a specific ‘MLLM’. We then aggregate results to thoroughly assess the performance of models and instructions using a variety of metrics. A colorful ‘Q’ symbolizes ‘Question’, and its design varies according to the specific task. For more detailed visual representations and explanations, please refer to Figure 5.
We introduce a comprehensive assessment framework named MM-InstructEval, which incorporates a diverse set of metrics to thoroughly evaluate various models and instructions within the realm of multimodal reasoning tasks with vison-text contexts, as depicted in Figure 2. MM-InstructEval is designed to complement existing zero-shot evaluation studies of Multimodal Large Language Models (MLLMs), offering a more comprehensive and holistic assessment in conjunction with prior research. Specifically, we introduce the “Best Performance” metric to explore how each model achieves its peak performance under specific instructions for each dataset. The metric establishes the upper benchmark for capabilities of each model on a given dataset. To assess the overall performance of each model across all instructions on a particular dataset, we propose the “Model Mean Relative Gain” metric. Similarly, the “Instruction Mean Relative Gain” metric is designed to evaluate the effectiveness of each instruction across all models. Stability is also a critical indicator of both model and instruction efficacy. A model that consistently delivers strong performance across various instructions is deemed more stable, while an instruction that consistently yields high performance across different models is considered highly effective. To measure robustness of models and instructions possessing superior performance, we introduce the “Model Stability” and “Instruction Stability” metrics. Furthermore, prior studies have typically focused either solely on assessing different models (Fu et al., 2023; Li et al., 2023a; Xu et al., 2023; Yu et al., 2023b; Yue et al., 2023; Ge et al., 2023) or exclusively on evaluating the performance of instructions (Chia et al., 2023; Ajith et al., 2023), often neglecting the crucial aspect of adaptability between models and instructions. To address this gap, we introduce the “Adaptability” metric to quantify how well different models adapt to various instructions. The metric evaluates how frequently each instruction achieves Top-K performance on a specific model across all datasets. Thereby, we can pinpoint which instructions consistently deliver superior performance for each model, effectively bridging the evaluation gap. Adaptability metric provides key insights into how models adjust to different instructional contexts, thereby enhancing the utility of MLLMs in practical applications. We conduct extensive experiments, and the results reveal significant insights into the performance of large (multimodal) language models in multimodal reasoning tasks with vision-text contexts: (1) Open-source vs. Closed-source: In most tasks, the performance of advanced closed-source models, such as GPT-4V and Claude3, generally exceeds that of all open-source models. Notably, these closed-source models demonstrate significantly superior performance in more challenging scenarios, such as the multimodal relation extraction task, etc. (2) Model Architecture and Performance: In the open-source group, models utilizing the Flan-T5 backbone (Chung et al., 2022), which is based on the Encoder-Decoder architecture, such as BLIP-2 (Li et al., 2023b) and InstructBLIP (Dai et al., 2023), outperform those in the LLaMA series (Touvron et al., 2023a, b), which employ a Decoder-only architecture. It suggests that the Encoder-Decoder framework may be more adept at handling the multimodal reasoning tasks possessing vison-text contexts. (3) Effectiveness of Instruction Formats: Instructions formatted as “Question-Answer” consistently yield better results. Furthermore, the inclusion of options within these instructions enhances performance on specific tasks, highlighting the importance of structured interaction in improving task comprehension and response accuracy. (4) Instruction Preferences to Models: Due to varied training paradigms, different models demonstrate distinct preferences for specific instructions. The variability underscores the need for adaptable and flexible instruction strategies to maximize model performance on diverse tasks. Our main contributions are summarized as follows:
  • We conduct zero-shot evaluations on 31 models across 16 datasets, utilizing 10 distinct instructions to cover 6 diverse multimodal reasoning tasks that depend on vision-text contexts. The comprehensive assessment, the first of its kind, evaluates a broad spectrum of MLLMs across various instructions in complex multimodal reasoning tasks. Our work builds upon and significantly extends previous evaluations of MLLMs, culminating in a thorough appraisal of cutting-edge MLLMs.
  • We introduce the MM-InstructEval framework, which incorporates a variety of metrics to provide a detailed evaluation of different models and instructions. Comprehensive metrics include the Best Performance metric to benchmark peak model capabilities, the Mean Relative Gain metric to assess overall efficacy, and the Stability metric to measure the robustness. Additionally, the Adaptability metric is proposed to quantify the adaptability between models and instructions, offering valuable insights for the design of instructions to maximize the performance of current and future models.
  • Through extensive experimentation, we establish benchmarks for both LLMs and MLLMs across various multimodal reasoning tasks with vision-text contexts. Our findings enrich the understanding of the interactions between model architectures, instruction formats, and task performance, providing strategic guidance for future advancements in applications of MLLMs.
2 RELATED WORK
2.1 LLMs-Based Evaluation
As Large Language Models (LLMs) (Yin et al., 2023) gain prominence, numerous evaluations of these models have surfaced. For example, Zhang et al. conduct comprehensive evaluations of LLMs, including ChatGPT and Flan-T5, across a variety of sentiment analysis tasks, ranging from text sentiment analysis to aspect-based sentiment analysis and multifaceted analysis of subjective texts. Additionally, Wang et al. assess the performance of ChatGPT across five representative sentiment analysis tasks. While most of the aforementioned studies primarily focus on text sentiment analysis and evaluate only a limited number of LLMs, there is a growing need for more comprehensive evaluation frameworks. To address this gap, LLMeBench introduces an open-source, user-friendly, and adaptable benchmarking framework designed specifically for LLMs. It features four key components: the Dataset module, Asset module, Model module, and Evaluation module. INSTRUCTEVAL (Chia et al., 2023) introduces a comprehensive evaluation suite specifically designed for 11 instruction-tuned large language models, including models such as Flan-T5, Vicuna, Alpaca, and others. In a related vein, InstructEval (Ajith et al., 2023) methodically explores the generalizability of popular instruction selection and induction methods for in-context learning (ICL) in LLMs.
While these studies are pivotal in evaluating LLMs, they predominantly focus on text-based tasks. Additionally, the aforementioned research tends to concentrate either solely on assessing different models or exclusively on evaluating the effectiveness of instructions, thereby overlooking the critical aspect of adaptability between models and instructions. Our primary focus is to assess the performance of various MLLMs across diverse multimodal reasoning tasks with vision-text contexts. To address the noted research gap, we propose the Adaptability metric to quantify the degree of adaptability between different models and various instructions, aiming to foster a deeper understanding of how MLLMs can be optimally utilized in complex, real-world scenarios.
2.2 MLLMs-Based Evaluation
Multimodal Large Language Models (MLLMs) (Yin et al., 2023), building upon the capabilities of LLMs, excel in a variety of multimodal tasks such as Caption Generation and Visual Question Answering. The proficiency spurs a surge in research of evaluating these models. We summarize the latest benchmarks for MLLMs in Table 1. Specifically, MME (Fu et al., 2023) introduces the first MLLM evaluation benchmark, categorizing tasks into Perception (Recognition, OCR, and others) and Cognition (Commonsense Reasoning, Numerical Calculation, Text Translation, and so on). MMBench (Liu et al., 2023c) provides a systematic framework to robustly evaluate the diverse capabilities of large vision-language models. SEED-Bench (Li et al., 2023a) evaluates comprehension across image and video modalities through 19K multiple-choice questions covering 12 dimensions. LVLM-eHub (Xu et al., 2023) assesses 8 MLLMs, including InstructBLIP and MiniGPT, on 47 visual text-related benchmarks, featuring quantitative evaluations and an online arena platform. MM-Vet (Yu et al., 2023b) benchmarks MLLMs across 16 tasks, identifying six core Visual Language (VL) capabilities essential for complex multimodal tasks. MMMU benchmark (Yue et al., 2023) aims to evaluate expert-level multimodal understanding capabilities of foundation models across a broad range of tasks. MLLM-Bench (Ge et al., 2023) utilizes small datasets and the capabilities of GPT-4V (OpenAI, 2023b) to assess the performance of various open-source MLLMs. Additional evaluations like HallusionBench (Guan et al., 2023) and Bingo (Cui et al., 2023) focus on the hallucination phenomenon in MLLMs.
Existing benchmarks primarily focus on vision-driven tasks where questions can typically be answered based solely on visual context. However, these benchmarks overlook multimodal reasoning tasks that require the integration of both visual and textual contexts to effectively answer questions. In such tasks, relying solely on visual content is insufficient for deriving correct answers, necessitating a deeper understanding and reasoning across modalities. We expand upon existing research by conducting a more extensive and comprehensive evaluation of MLLMs in diverse reasoning tasks involving complex multimodal contexts. Our evaluation not only enhances the understanding about capabilities of MLLMs in complex multimodal environments but also advances knowledge in the field.
Figure 3: Inference process of Multimodal Language Models (MLLMs) for AlgoPuzzleVQA employing varied multimodal instructions. We construct instructions based on these formats, encompassing mandatory components, such as Task name, Task definition, and Output format, Question, as well as optional components, for instance, Context and Options. Furthermore, each format incorporates Specific instruction trigger words customized for the respective instruction. Note that only text context are provided for inputting the Large Language Models (LLMs).
Figure 4: Inference process of Multimodal Language Models (MLLMs) for MSD employing varied multimodal instructions.
3 MM-InstructEval
We introduce MM-InstructEval, an extensive evaluation framework that integrates a diverse set of metrics to thoroughly assess various models and instructions in multimodal reasoning tasks involving vision-text contexts, as depicted in Figure 2. In the following sections, we will delve into a detailed exploration of the definition of multimodal reasoning tasks that incorporate multimodal contexts. Additionally, we will outline the design of various instructions and discuss the metrics proposed for evaluating the performance of models and instructions.
3.1 Definition of Multimodal Reasoning Tasks with Vision-Text Contexts
Prior research on multimodal evaluation, highlighted in studies such as (Fu et al., 2023; Li et al., 2023a; Xu et al., 2023; Yu et al., 2023b), often confines the evaluation of Multimodal Large Language Models (MLLMs) to a basic assessment of their multimodal capabilities. These studies typically examine the interplay between textual instructions/questions and vision-only contexts, primarily testing traditional vision-language (VL) capabilities (Yu et al., 2023b), which encompass Language Generation, Recognition, Knowledge Reasoning, Spatial Reasoning, OCR, and more. A significant limitation of such evaluations is their inadequacy in fully assessing the capabilities of MLLMs to handle multimodal reasoning tasks that require combined information from both vision and text contexts. For example, in tasks like Multimodal Sarcasm Detection (MSD), shown in Figure 1, detecting sarcasm through visual context alone is insufficient; accurate detection necessitates the integration of both vision and text contexts. It is crucial to recognize that tasks involving complex multimodal contexts demand substantial multimodal reasoning abilities. These tasks require extracting information from text, aligning it with visual data, and engaging in reasoning processes that inherently necessitate multimodal thinking.
Inspired by MM-Vet (Yu et al., 2023b), we delineate conventional vision-language (VL) capabilities for evaluation, with examples illustrated in the upper part of Figure 1. In contrast, our research delves into multimodal reasoning tasks that incorporate vision-text contexts. These tasks require not just interactions between textual instructions and visual contexts to harness vision-language multimodal capabilities as described previously, but also an in-depth understanding of multimodal contexts, as demonstrated in Figure 1. Illustrative examples are provided in Figure 3 and Figure 4, showcasing how both textual content and images are crucial for formulating responses to the given instructions. We aim to conduct a comprehensive evaluation of various MLLMs, thoroughly covering a range of multimodal reasoning tasks that integrate vision-text contexts, encompassing Visual Question Answering with Multimodal Contexts (VQAMC), Multimodal Sentiment Analysis (MSA), Multimodal Aspect-Based Sentiment Analysis (MABSA), Multimodal Hateful Memes Detection (MHMD), Multimodal Sarcasm Detection (MSD), and Multimodal Relation Extraction (MRE). To our knowledge, it is the first work to evaluate such a diverse array of tasks using MLLMs. Notably, apart from VQAMC, existing benchmarks have not extensively explored the performance of MLLMs across various multimodal reasoning tasks possessing compelx multimodal contexts. Ultimately, we delineate the essential Visual Language (VL) capabilities necessary for evaluating multimodal reasoning tasks with vision-text contexts, including Multimodal Alignment, Multimodal Fusion, Multimodal Content Understanding, and Multimodal Reasoning. Illustrative examples of these capabilities are showcased in the lower part of Figure 1.
  • $\bullet$
    Multimodal Alignment: It refers to the process of associating or aligning information from different modalities in a coherent and meaningful manner. The objective is to establish correspondences or connections between different modalities to enable models to effectively leverage information from diverse sources. For instance, in multimodal models processing both images and text, alignment might involve matching specific words or phrases in the text with corresponding objects or concepts in the images. An example is aligning the phrase “Lady Gaga” with the “Person region” on the left side of the image in the “MABSA” column of Figure 1. Such alignment fosters a more comprehensive and integrated understanding of the content, empowering the model to generate more contextually relevant responses or undertake tasks requiring a joint comprehension of multiple modalities.
  • $\bullet$
    Multimodal Fusion: The process involves integrating information from multiple modalities to achieve a more comprehensive representation, thereby enhancing the overall understanding of multimodal data and improving model performance. For example, as depicted in Figure 1, the fusion of the phrase “summer fun day” with the visual context of “the blue sunny sky” in the image within the “MSA” column amplifies the positive sentiment conveyed by the user. The fusion not only enriches the interpretation of individual modal inputs but also enables a more precise response from the model.
  • $\bullet$
    Multimodal Content Understanding: The capability involves the ability of a model to comprehend and interpret information presented across various modalities, such as the text and vision. For instance, in the “MSD” column of Figure 1, there is a juxtaposition between the text context and the vision context, which collectively convey the irony of user. The scenario highlights the necessity for high-level understanding capabilities of the model to accurately interpret and respond to the complex, layered meanings embedded in multimodal inputs.
  • $\bullet$
    Multimodal Reasoning: The process involves drawing logical conclusions or making predictions based on available information or evidence. In tasks that utilize multimodal contexts, the expected output may not be explicitly present in the original content. It requires the model to engage in reasoning based on the combined image and text contexts to derive the appropriate response. Such tasks demand inference from the integrated vision-text context to formulate the final answer, highlighting the need for sophisticated reasoning capabilities in handling complex multimodal data.
Figure 5: Details of the various components of multimodal instructions for different tasks, such as Multimodal Sentiment Analysis (MSA), Multimodal Aspect-Based Sentiment Analysis (MABSA), Multimodal Hateful Memes Detection (MHMD), Multimodal Sarcasm Detection (MSD), and Multimodal Relation Extraction (MRE).
3.2 Various Multimodal Instructions
Recent research in Natural Language Processing (NLP) has demonstrated that variations in instructions, even with identical semantics, can significantly affect the performance of a model (Perez et al., 2021; Lu et al., 2022b; Zhang et al., 2023b; Liu et al., 2023b; Gan and Mori, 2023). Building on the insights from Zhang et al., we craft multimodal instructions for diverse multimodal reasoning tasks incorporating vision-text contexts, which include essential elements such as the task name, definition, and output format, as well as optional components like answer options. The text and image contexts form the multimodal content for each instance across various multimodal reasoning tasks. These instructions are specifically designed to evaluate how different MLLMs perform in a zero-shot setting when faced with diverse instructions. Although we develop only 10 manual instructions to gauge the performance of various models across different datasets, the evaluations offer valuable insights, enabling us to draw significant conclusions. For instance, we design a range of multimodal instructions for the Visual Question Answering with Multimodal Contexts task, as depicted in Figure 3. Similarly, consistent instruction structures that integrate both text and image contexts also are created for other tasks. The instructions tailored for the Multimodal Sarcasm Detection (MSD) task are detailed in Figure 4.
Specifically, the Task name component serves to identify and clarify the purpose of each multimodal task. The Task definition component, informed by the definition of task and annotation guidelines, outlines the label space as a set of options from which the model can generate responses. The Output format component dictates the expected structure of the output of model, aiding in decoding the response into the desired format. The terms Question and Text context are essential in multimodal reasoning tasks involving vision-text contexts. The Options component is optional and provides the model with multiple-choice questions, guiding it to respond accordingly. We observe that various MLLMs exhibit preferences for specific instructions. The preference may be influenced by the use of Specific instruction trigger words during training of models, such as “### Instruction:” or “Question:”. Details of these components in text instructions for various tasks are illustrated in Figure 5.
3.3 Comprehensive Metrics
To effectively evaluate the performance of different models across various tasks, we develop a suite of comprehensive metrics. These metrics include the Best Performance metric, the Mean Relative Gain metric, the Stability metric, and the Adaptability metric. Each is designed to assess the efficacy of different models ($\mathcal{M}$) and various instructions ($\mathcal{I}$) across multiple datasets ($\mathcal{D}$), covering a range of diverse multimodal reasoning tasks with vison-text contexts.
We choose Accuracy, denoted as accmdi, as primary metric to assess the performance of a model, $m\in\mathcal{M}$, with a specific multimodal instruction, $i\in\mathcal{I}$, on each dataset, $d\in\mathcal{D}$.
$p^{j}_{mdi}=LM(T^{j},V^{j}),\ acc_{mdi}=\frac{\sum_{j=1}^{N_{d}}{\mathbbm{1}}_{(p^{j}_{mdi}=l^{j}_{mdi})}}{N_{d}},$ (1)
where LM represents a specific language model, the LLM takes only the text instruction, including text context, as input, denoted as $T$, while the MLLM takes the multimodal context, $(T,V)$, as input. For the $j$-th instance, $p^{j}_{mdi}$ represents the predicted label of the LM, $l^{j}_{mdi}$ is the true label. $\mathbbm{1}$ means the indicator function, $N_{d}$ is the number of instances, $acc_{mdi}\in Acc_{mdi}$ is the accuracy performance, $Acc_{mdi}\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{D}|\times|\mathcal{I}|}$, |||*|| ∗ | means the number of *. $|\mathcal{M}|=31$, $|\mathcal{D}|=16$, and $|\mathcal{I}|=10$ represent the number of evaluated models, datasets, and instructions, respectively.
Best Performance.
To account for performance variations across different instructions, we report the best accuracy, denoted as $A^{\tilde{i}}\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{D}|}$, which represents the highest accuracy achieved by each model among all instructions on each dataset. The metric highlights the upper performance limits for different models, providing insight into their optimal capabilities under various instructional scenarios. We acquire the best accuracy $A^{\tilde{i}}_{md}$.
$A_{md}$ $=Max(\{(acc_{mdi})\}_{i\in\mathcal{I}}),$ (2)
$\tilde{i}_{md}$ $=\mathop{\arg\max}_{i}\ \ \ (\{(acc_{mdi})\}_{i\in\mathcal{I}}),$
Mean Relative Gain (MRG).
[…]
$MRG^{\mathcal{M}}_{md}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}r_{mdi},MRG^{\mathcal{M}}\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{D}|},$ (3)
$r_{mdi}=\frac{acc_{mdi}-\overline{Acc_{di}}}{\overline{Acc_{di}}}\times 100,$
$\overline{Acc_{di}}=\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}Acc_{mdi},\overline{Acc_{di}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{I}|}.$
[…]
$MRG_{id}^{\mathcal{I}}=\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}r_{idm},MRG^{\mathcal{I}}\in\mathbb{R}^{|\mathcal{I}|\times|\mathcal{D}|},$ (4)
$r_{idm}=\frac{acc_{idm}-\overline{Acc_{dm}}}{\overline{Acc_{dm}}}\times 100,$
$\overline{Acc_{dm}}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}Acc_{idm},\overline{Acc_{dm}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{M}|},$
where the $Acc_{idm}$ is the transpose of $Acc_{mdi}$.
Stability.
Stability is a crucial indicator of the reliability of both models and instructions. We assess the stability of a model across various instructions using the Model Stability Metric, denoted as ${}^{\mathcal{M^{{}^{\prime}}}}$. The metric is derived by calculating the standard deviation of the accuracy for each model when the instructions are varied for each dataset. A lower standard deviation indicates a model that performs consistently across different instructional contexts.
$S^{\mathcal{M}^{{}^{\prime}}}_{md}=\sqrt{\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}(acc_{mdi}-\overline{Acc_{md}})^{2}}.$ (5)
Similarly, the stability of an instruction across different models is measured using the Instruction Stability Metric, represented as ${}^{\mathcal{I^{{}^{\prime}}}}$. The metric is calculated by computing the standard deviation of the accuracy for each instruction across different models for each dataset. Like ${}^{\mathcal{M^{{}^{\prime}}}}$, a lower standard deviation here signifies that the instruction performs consistently across various models.
$S^{\mathcal{I}^{{}^{\prime}}}_{id}=\sqrt{\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}(acc_{idm}-\overline{Acc_{id}})^{2}},$ (6)
where $\overline{Acc_{md}}$ and $\overline{Acc_{id}}$ are the transpose of $\overline{Acc_{dm}}$ and $\overline{Acc_{di}}$, respectively. $\mathcal{M}^{{}^{\prime}}\subset\mathcal{M}$, and $\mathcal{I}^{{}^{\prime}}\subset\mathcal{I}$.
Adaptability.
Different instructions significantly influence model performance. To quantify the adaptability between models and instructions, we introduce the Global Top-K Hit Ratio ($GHR@K$) as a metric to evaluate how each model performs for various instructions across all datasets. The metric measures the frequency at which each instruction achieves Top-K performance on a specific model across all datasets. Our goal is to pinpoint specific instructions that enable each model to consistently achieve high performance across all datasets. Consequently, we assess each instruction based on its overall Top-K Hit Ratio score across all datasets for model $m$, providing a clear picture of which instructions are most effective universally.
$I^{K}_{d|m}=\mathop{\arg\max}_{i^{1},...,i^{K}}K\ \ \ (\{(acc_{mdi})\}_{i\in\mathcal{I}}),|I^{K}_{d|m}|=K,$ (7)
$GHR@K_{m}=\frac{Counter_{m}^{\mathcal{I}}([I^{K}_{d^{1}|m}||...||I^{K}_{d^{|\mathcal{D}|}|m}])}{(|\mathcal{D}|\times K)},$
[…]
4 Comprehensive Evaluation
In this section, we offer a comprehensive overview of the models evaluated by MM-InstructEval, as well as the various multimodal reasoning tasks with vision-text contexts and their corresponding datasets. Specifically, we assess a range of language models (LMs) across different multimodal reasoning tasks, as detailed in Table 3. It includes both pure Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs). We consider both closed and open-source models within each category, aiming to include as many currently popular models as possible. Detailed information about these models is provided in Table 2. MM-InstructEval is designed to be flexible, easily accommodating emerging models and extendable to new datasets. We continuously update and enrich it. Detailed and latest results can be found on the MM-InstructEval Leaderboard. Our evaluations are performed in a zero-shot setting, aiming to assess the capabilities of models in multimodal reasoning tasks with multimodal contexts, enabling them to reason accurate answers without the need for fine-tuning or few-shot demonstrations on our benchmarks. All experiments [^1] Different seeds have a limited impact on the experimental results. We only use seed 42 for experiments, which allows us to still draw valuable conclusions. are performed using NVIDIA A6000 GPUs.
4.1 Why do Zero-Shot Evaluation?
Following previous research (Fu et al., 2023; Liu et al., 2023c; Li et al., 2023a; Xu et al., 2023; Yu et al., 2023b; Yue et al., 2023; Ge et al., 2023) that evaluates MLLMs, we concentrate on assessing zero-shot performance across a variety of multimodal reasoning tasks with vision-text contexts for the following reasons:
  • Zero-shot evaluation serves as a fundamental and rigorous test, measuring the intrinsic ability of models to handle diverse multimodal reasoning tasks with complex contexts. It allows us to draw valuable conclusions about the capabilities of models without prior specialized training.
  • Most MLLMs are designed to process single image-text pairs, and the format of “multiple image-text pairs” does not align with their training configurations, indicating limited compatibility with few-shot evaluation settings.
  • Preliminary experiments to assess the in-context capabilities of a few MLLMs yielded suboptimal results. Looking ahead, we plan to develop a tailored subset of data and design effective methods to facilitate in-context evaluation, aiming to enhance the performance of MLLMs in such scenarios.
4.2 Evaluated Models
Below is a detailed overview of the various language models employed in our evaluation.
Table 2: Comprehensive summary of various models (Red indicates closed-source models; others are open-source). Below are the abbreviations used in the table: ‘LLMs’ refers to the pretrained Language Models (LLMs) that form the backbone of the Multimodal Large Language Models (MLLMs). ‘PVM’ signifies the pretrained visual model backbone of the MLLMs. ‘To-Paras’ and ‘Tr-Paras’ represent the total number of parameters and trainable parameters for each language model, respectively. ‘Held-In’ refers to the dataset that each specific MLLM was trained or fine-tuned on. The ‘GPU’ column indicates the single GPU utilization during inference, and the ‘Time’ column signifies the time taken for model inference on each text/multimodal instance using a single GPU. Note that GPU usage and inference time may vary slightly across different datasets due to varying data lengths. For these two metrics, we provide an approximate mean value across all datasets for each model. A dash, ‘-’, indicates that the data is not applicable or not involved, and ‘Unknown’ means we do not have information on whether these models cover the evaluated dataset.
Modality Models LLMs PVMs To-Paras Tr-Paras Held-In GPU Time
Text ChatGPT gpt-3.5-turb - - - Unknown - -
LLaMA1-7B LLaMA-V1-7B - 6.74B 6.74B - 26G 2.0s
LLaMA1-13B LLaMA-V1-13B - 13.02B 13.02B - 48G 9.0s
LLaMA2-7B LLaMA-V2-7B - 6.74B 6.74B - 26G 1.0s
LLaMA2-13B LLaMA-V2-13B - 13.02B 13.02B - 48G 8.0s
Mixtral-AWQ Mixtral-8x7B-Instruct-v0.1 - 0.26B 0.26B Unknown 25G 5s
Gemma Gemma-7B-it - 8.54B 8.54B Unknown 36G 1.5s
Flan-T5-XXL Flan-T5-XXL - 11.14B 11.14B - 44G 0.3s
Multimodal GPT-4V gpt-4 vision-preview - - Unknown - -
Gemini-V gemini-pro-vision - - - Unknown - -
Claude3 claude-3-opus-$2$ Unknown - -
OpenFlamingo LLaMA-7B ViT-L/14 8.34B 1.31B 34G 1.5s
Fromage OPT-6.7B ViT-L/14 6.97B 0.21B - 14G 5.0s
LLaVA-v0-7B Vicuna-7B-v0 ViT-L/14 6.74B 6.74B - 15G 2.5s
LLaVA-v0-13B Vicuna-13B-v0 ViT-L/14 13.02B 13.02B - 27G 2.0s
LLaVA-v1.6-7B Vicuna-7B ViT-L/14 7.06B 6.76B - 26G 1.3s
LLaVA-v1.6-13B Vicuna-13B ViT-L/14 13.35B 13.05B - 36G 1.3s
MiniGPT4 Vicuna-13B ViT-g/14 14.11B 0.04B - 15G 1.3s
mPLUG-Owl LLaMA-V1-7B ViT-L/14 7.12B 7.12B - 16G 1.0s
mPLUG-Owl2.1 Qwen-7B ViT-G/14 11.42B 11.42B - 24G 1.0s
LLaMA-Adapter V2 LLaMA-V1-7B ViT-L/14 7.23B 7.23B - 14G 1.3s
VPGTrans Vicuna-7B - 7.83B 0.11B - 36G 10s
Multimodal-GPT LLaMA-V1-7B ViT-L-14 8.37B 0.02B - 18G 0.5s
LaVIN-7B LLaMA-V1-7B ViT-L/14 7.17B 7.17B ScienceQA 16G 4.0s
LaVIN-13B LLaMA-V1-13B ViT-L/14 13.36B 13.36B ScienceQA 28G 11.0s
Lynx Vicuna-7B Eva-ViT-1b 8.41B 0.69B Hate 44G 6.5s
Fuyu-8B Persimmon-8B - 9.41B 9.41B Unknown 44G 1.3s
LaVIT LLaMA-V2-7B ViT-G/14 8.30B 0.41B - 23G 2.2s
Qwen-VL-Chat Qwen-7B ViT-bigG 9.66B 9.66B - 22G 1.0s
Multimodal BLIP-2 Flan-T5-XXL ViT-G/14 12.23B 0.11B - 26G 3.5s
InstructBLIP Flan-T5-XXL ViT-G/14 12.31B 0.45B - 16G 0.3s
4.2.1 LLMs
We evaluate eight LLMs (closed-source and open-source) across multiple tasks using text-only contexts. Since the questions from the AlgoPuzzleVQA and MMMU datasets require visual content for answers, we do not include these two datasets in our evaluation of LLMs.
  • ChatGPT (OpenAI, 2023a) is a closed-source conversational AI language model developed by OpenAI, renowned for its impressive performance across a wide range of NLP tasks. Specifically, we evaluate the classic model, “ChatGPT (gpt-3.5-turbo)” [^2] Our evaluation of ChatGPT is carried out between July and September 2023. .
  • The LLaMA family of models includes LLaMA-1 (Touvron et al., 2023a) and LLaMA-2 (Touvron et al., 2023b). We assess the performance of the “decapoda-llama-7b/13b-hf” and “meta-Llama-2-7b/13b-hf” models.
  • Mixtral (Mixtral-AI, 2023.11) is a sparse mixture-of-experts network (SMoE). Due to its high GPU resource requirements, which exceed our server capabilities, we employ a version with weight quantization using AWQ [^3] AWQ is an efficient, accurate, and extremely fast low-bit weight quantization method that supports 4-bit quantization.. Specifically, we evaluate the “ybelkada/Mixtral-8x7B-Instruct-v0.1-AWQ” model in our study.
  • Gemma (Google, 2024.02) represents a new family of state-of-the-art open-source LLMs by Google, developed under the Gemini initiative [^4] https://deepmind.google/technologies/gemini/#gemini-1.5. In our study, we evaluate the “google/gemma-7b-it” model.
  • Flan-T5 (Chung et al., 2022) is a model that extends its capabilities to 1836 fine-tuning tasks through instruction tuning, enhancing the performance and usability of the model. In our evaluation, we specifically examine the “flan-t5-xxl” version.
4.2.2 MLLMs
We evaluate a total of 23 Multimodal Large Language Models. The closed-source models include GPT-4V (OpenAI, 2023b), Gemini-V (Google, 2023), and Claude3 (Anthropic, 2024). The open-source MLLMs are developed by integrating pretrained visual models (PVMs) with LLM backbones. Typically, the PVMs are derived from pretrained CLIP models, such as ViT-L/14, ViT-g/14, etc. The LLMs span various families including the LLaMA family, with models like LLaMA-V1 and LLaMA-V2, Vicuna, the Flan-T5 family, and others.
[…]
  • Claude3-V represents the latest in cutting-edge closed-source models. Claude (Anthropic, 2024), developed by Anthropic, is a family of large language models designed to transform interactions with AI. For our evaluation, we utilize the “claude-3-opus-20240229” [^6] Our evaluation of Claude3-V via the corresponding API is carried out in March 2024. version.
  • Gemini-V is capable of processing both text and image inputs to generate text responses. Gemini (Google, 2023), developed by Google, is a series of advanced multimodal generative AI models. In our study, we evaluate the “gemini-pro-vision” [^7] Our evaluation of Gemini-V via the corresponding API was conducted in January 2024. version.
  • The OpenFlamingo family of models is renowned for its impressive performance in few-shot learning across various open-ended vision and language tasks, as detailed in (Alayrac et al., 2022; Awadalla et al., 2023). We evaluate the pre-trained “Openflamingo-9B” model.
  • Fromage is trained by visually grounding LLMs through image captioning and contrastive learning (Koh et al., 2023). We assess the pre-trained “fromage-model”.
  • LLaVA (Liu et al., 2023a) is an end-to-end traine MLLM that connects a vision encoder with an LLM for extensive visual and language understanding. We assess different pretrained LLaVA models, which are built on various scales of LLMs, including “LLaVA-v0-7B”, “LLaVA-v0-13B”, “LLaVA-v1.6-7B”, and “LLaVA-v1.6-13B”.
  • MiniGPT-4 aligns a frozen visual encoder with a frozen LLM called Vicuna, using one projection layer (Zhu et al., 2023). We utilize the pretrained “MiniGPT-4 checkpoint with Vicuna 13B”.
  • mPLUG-Owl equips LLMs with multi-modal abilities through modularized learning of a foundation LLM, a visual knowledge module, and a visual abstractor module (Ye et al., 2023a, b). We evaluate the “mPLUG-Owl 7B” and “mPLUG-Qwl2.1”.
  • LLaMA-Adapter V2 (Gao et al., 2023) is jointly trained on image-text pairs and instruction-following data. We evaluate the “LLaMA-Adapter V2 Multimodal” model.
  • VPGTrans (Zhang et al., 2023a) is a two-stage transfer framework facilitates the efficient transfer of Visual Prompt Generators (VPG) across LLMs while minimizing the need for extensive training data. We assess the performance of the “VL-LLaMA” model, in which the VPG is seamlessly transferred from BLIP-2 OPT6.7B to LLaMA-7B.
  • Multimodal-GPT (Gong et al., 2023) incorporates the Low-rank Adapter (LoRA) (Hu et al., 2022) in both the gated-cross-attention and self-attention components of the language model. We assess the pretrained “mmgpt-lora-v0-release” weights.
  • LaVIN (Luo et al., 2023) is a model proposed based on the concept of Mixture-of-Modality Adaptation (MMA). We conduct evaluation of the “LaVIN-7B” and “LaVIN-13B”.
  • Lynx (Zeng et al., 2023) is a straightforward prefix-tuning GPT4-style model featuring a two-stage training approach. We utilize the “finetune-lynx.pt” model for evaluation.
  • Fuyu-8B (AI, 2023) is a multimodal text and image converter trained by Adept AI [^8] https://www.adept.ai/. Architecturally, Fuyu is a pure decoder transformer - there is no image encoder and instead utilizes a transformer decoder. “adept/Fuyu-8B” is assessed.
  • LaVIT (Jin et al., 2023) is a new effective, general-purpose multimodal foundation model that goes beyond traditional adapter-based architectures. “rain1011/LaVIT-7B-v2” is evaluated in our paper.
  • Qwen-VLs (Bai et al., 2023) are a series of highly performant and versatile vision-language foundation models based on Qwen-7B (Qwe, 2023) language model by introducing a new visual receptor. We evaluate the “Qwen-VL-Chat”.
  • BLIP-2 (Li et al., 2023b) addresses the modality gap through the use of a lightweight Querying Transformer that connects a frozen pre-trained image model with a language model. We consider the version of this model labeled as “blip2-flan-t5-xxl”.
  • InstructBLIP (Dai et al., 2023) conducts a comprehensive and systematic study on vision-language instruction tuning by utilizing pretrained BLIP-2 models. We utilize the pretrained model “blip2-instruct-Flan-T5xxl”.
4.3 Multimodal Datasets
Table 3: Evaluation datasets statistics for different tasks. “Labels” denotes the number of labels among each task. “-” means that the VQAMC task does not have a fixed label space, and each question has a unique answer.
Task Dataset Modality Test Labels Metric
VQAMC ScienceQA Text-Image 2017 - Acc
AlgoPuzzleVQA Text-Image 1800 - Acc
MMMU-val Text-Image 900 - Acc
MSA MVSA-Single Text-Image 413 3 Acc
MVSA-Multiple Text-Image 1531 3 Acc
TumEmo Text-Image 9463 7 Acc
MOSI-2 Video 654 2 Acc
MOSI-7 Video 684 7 Acc
MOSEI-2 Video 2797 2 Acc
MOSEI-7 Video 3588 7 Acc
MABSA Twitter-2015 Text-Image 1037 3 Acc
Twitter-2017 Text-Image 1234 3 Acc
MASAD Text-Image 4935 2 Acc
MHMD Hate Text-Image 500 2 Acc
MSD Sarcasm Text-Image 2409 2 Acc
MRE MNRE Text-Image 640 19 Acc
We conduct comprehensive evaluation of various Language Models across a diverse range of multimodal reasoning tasks, including VQAMC, MSA, MABSA, MHMD, MSD, and MRE. Detailed statistics for each task and the datasets can be found in Table 3. The detailed introduction to various multimodal reasoning tasks with vision-text contexts and related datasets are as follows.
  • Visual Question Answering with Multimodal Contexts (VQAMC): ScienceQA (Lu et al., 2022a) is a well-known visual question-answering dataset featuring a variety of science topics. It provides both image-text contexts, which can range from semantically rich information to simple hints. Given our evaluation, we specifically utilize the portion of the dataset that includes image context. AlgoPuzzleVQA (Ghosal et al., 2024) is a novel dataset designed to challenge and evaluate the capabilities of MLLMs in solving algorithmic puzzles that require visual understanding, language comprehension, and complex algorithmic reasoning. The dataset includes areas such as boolean logic, combinatorics, graph theory, optimization, and search. It is crucial to note that the AlgoPuzzleVQA dataset requires the integration of textual context with visual elements to resolve specific queries effectively. Therefore, our assessment focuses exclusively on MLLMs capable of handling such intricate, dual-modality data interpretation. For simplicity, we refer to this dataset as “PuzzleVQA”. MMMU (Yue et al., 2023) presents significant challenges as an extensive benchmark designed for college-level multi-discipline multimodal understanding and reasoning. It encompasses six common disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering. Our evaluations specifically focus on the validation set of MMMU, succinctly referred to as “MMMU-val.”
  • Multimodal Sentiment Analysis (MSA): MSA aims to detect the overall sentiment of a text-image pair or a video, addressing the complexity of sentiments expressed across different modalities (Xu et al., 2018; Yang et al., 2021a, b; Li et al., 2022; Yang et al., 2023a). Our evaluation includes three widely used text-image datasets: MVSA-Single and MVSA-Multiple (Niu et al., 2016), and TumEmo (Yang et al., 2021a). Additionally, we assess four video datasets: MOSI-2, MOSI-7 (Zadeh et al., 2016), MOSEI-2, and MOSEI-7 (Zadeh et al., 2018). Given that most MLLMs primarily accept text-image pairs, for video datasets, we initially extract one frame per second to form a set of candidate frames. We then randomly select a single frame [^9] We also experiment with selecting multiple frames, such as three, and inputting them individually into the MLLMs, using a voting principle to determine the final result. However, results are comparable to using a single randomly chosen frame. Thus, we opt for one frame as the visual input for each video. to serve as the image input for MLLMs. For MOSI-2 and MOSEI-2, the label space includes {positive, negative}, excluding neutral, which is labeled as zero. In contrast, MOSI-7 and MOSEI-7 feature a label space comprising {strongly positive, positive, weakly positive, neutral, weakly negative, negative, strongly negative}, with neutral sentiment explicitly recognized.
  • Multimodal Aspect-Based Sentiment Analysis (MABSA): MABSA focuses on identifying sentiments associated with specific aspect terms within the given vision-text contexts (Hu et al., 2019; Yu and Jiang, 2019; Ju et al., 2021; Ling et al., 2022; Yang et al., 2022a; Yu et al., 2023a; Yang et al., 2022b; Yu et al., 2022). Our experiments utilize three widely recognized datasets: Twitter-2015 (Zhang et al., 2018), Twitter-2017 (Lu et al., 2018), and MASAD (Zhou et al., 2021). These datasets evaluate the effectiveness of MLLMs in accurately discerning sentiments related to specific aspects based on diverse multimodal contexts.
  • Multimodal Hateful Memes Detection (MHMD): A new challenge set for multimodal classification, specifically designed for MHMD, is introduced in (Mathias et al., 2021). However, a labeled test set has not been publicly released for this challenge. Consequently, our evaluation focuses on the publicly available labeled validation set, commonly referred to as “dev-seen” in the literature. For simplicity, we refer to the dataset as Hate in this paper.
  • Multimodal Sarcasm Detection (MSD): MSD aims to identify sarcasm with multimodal contents, drawing upon the interplay of text and visual cues (Cai et al., 2019; Liang et al., 2022; Qiao et al., 2023). Cai et al. introduce a new dataset based on Twitter specifically tailored for multimodal sarcasm detection. In our paper, we refer to this dataset as Sarcasm.
  • Multimodal Relation Extraction (MRE): MRE involves identifying relations between two entities, supported by visual-text contexts (Zheng et al., 2021; Xu et al., 2022; Yuan et al., 2023). To support research in this domain, Zheng et al. introduce the Multimodal Neural Relation Extraction dataset (MNRE), which is manually labeled and serves as a valuable resource for MRE tasks.
Table 4: The best zero-shot performance, $A^{\tilde{i}}$ ($\uparrow$, measured by Accuracy), of various large language models on different datasets. Superscripts indicates the specific instruction that yields the best performance for each dataset on the respective model. The “Total” column represents the sum of accuracy scores across all datasets for each model. MVSA-S, MVSA-M, Twitter15, and Twitter17 refer to the MVSA-Single, MVSA-Multiple, Twitter-2015, and Twitter-2017 datasets, respectively. AdapterV2 and MultiGPT represent the LLaMA-AdapterV2 and Multimodal-GPT models, respectively. $\heartsuit$ indicates that the specific closed-source models (GPT-4V and Claude3-V) are evaluated on subsets. The highlighted text indicates the best performance on the full datasets. The highlighted text denotes superior performance on the subsets, outperforming the top results of the model evaluated across the full dataset. The notation “-” signifies that we do not assess LLMs for the PuzzleVQA and MMMU datasets. The exclusion is due to the inherent requirement of these two datasets for both image information and textual context to generate effective responses. The “Total” results of LLMs are also not calculated. The “Total” represents the aggregate results across all datasets, with the exception of PuzzleVQA and MMMU.
Models VQAMC MSA MABSA MHMD MSD MRE Total Total
ScienceQA PuzzleVQA MMMU MVSA-S MVSA-M TumEmo MOSI-2 MOSI-7 MOSEI-2 MOSEI-7 Twitter15 Twitter17 MASAD Hate Sarcasm MNRE
ChatGPT 69.417 - - 56.553 53.187 48.173 89.605 44.4410 84.975 40.771 65.484 59.9710 72.703 60.848 69.027 38.282 - 853.38
LLaMA1-7B 36.195 - - 67.231 60.721 38.261 82.012 34.262 75.621 15.501 58.533 46.433 65.677 50.404 58.994 2.663 - 692.47
LLaMA1-13B 43.336 - - 66.992 68.828 44.686 72.105 34.112 79.552 28.742 52.073 47.247 65.492 49.205 57.534 19.225 - 729.07
LLaMA2-7B 43.086 - - 66.991 69.222 40.284 67.681 26.381 77.301 16.781 58.533 46.602 67.192 52.001 56.331 3.597 - 691.95
LLaMA2-13B 55.787 - - 66.021 68.697 45.786 81.862 31.496 81.662 24.336 60.375 48.545 69.102 55.002 60.231 20.005 - 768.85
Mixtral-AWQ 66.633 - - 54.379 55.5910 43.941 87.925 41.237 79.305 38.277 55.451 60.212 70.561 58.005 64.387 37.039 - 812.88
Gemma 52.955 - - 67.728 61.619 43.102 81.657 36.117 77.052 25.067 54.295 52.435 70.945 57.801 60.072 30.312 - 771.09
Flan-T5-XXL 67.439 - - 64.814 66.014 49.563 89.6010 42.866 86.526 46.296 72.135 63.703 74.398 57.407 71.409 31.419 - 883.51
GPT-4V 79.8010 31.113 57.789 76.195 71.054 50.583 90.913 61.193 87.108 49.441 53.858 60.162 78.953 78.008 76.769 54.298 1057.16 968.27
Claude3-V 76.355 36.1110 53.337 80.954 69.085 46.1510 78.791 32.841 79.937 30.731 38.465 54.475 76.321 66.005 71.372 57.1410 948.02 858.58
Gemini-V 80.355 26.727 47.117 72.733 70.183 51.653 88.348 48.688 87.147 38.867 54.518 59.328 81.023 68.324 56.8310 41.414 973.17 899.34
OpenFlamingo 39.275 23.675 24.447 55.587 61.157 29.479 79.977 24.852 77.307 12.122 57.285 46.195 66.917 49.402 52.681 3.126 703.40 655.29
Fromage 34.517 21.947 20.337 29.856 28.192 22.761 57.197 19.152 47.412 11.042 19.966 27.316 35.106 37.602 40.687 0.161 453.18 410.91
LLaVA-v0-7B 41.105 19.945 24.854 69.423 65.426 30.448 74.692 30.039 74.657 18.129 35.103 44.573 73.483 39.805 43.2110 4.068 688.88 644.09
LLaVA-v0-13B 47.747 12.612 15.187 73.067 69.612 38.514 80.187 30.906 76.587 28.373 37.995 48.465 77.699 40.405 44.295 3.915 725.48 697.69
LLaVA-v1.6-7B 65.895 28.114 34.767 59.952 67.232 45.121 85.639 30.416 81.621 23.977 59.319 52.841 75.189 57.802 59.615 17.191 844.62 781.75
LLaVA-v1.6-13B 69.665 26.899 39.533 64.563 60.433 53.223 86.393 35.969 78.263 40.897 58.737 56.0810 77.9510 61.204 62.313 22.813 894.87 828.45
MiniGPT4 58.705 27.335 30.565 71.125 70.782 50.294 83.994 35.422 83.382 38.465 47.167 49.439 76.012 54.2010 57.493 2.816 837.13 779.24
mPLUG-Owl 37.937 26.173 26.226 51.946 50.366 33.371 68.751 28.286 58.106 20.296 33.752 38.742 58.262 48.601 49.737 5.475 635.96 583.57
mPLUG-Owl2.1 60.295 25.611 33.332 53.647 63.117 47.025 85.635 35.675 73.477 33.926 60.665 55.115 75.605 58.205 60.485 22.193 843.93 784.99
AdapterV2 54.447 25.727 27.113 73.547 70.134 39.1410 86.438 38.348 82.028 33.539 37.329 48.389 76.294 50.407 57.205 7.193 807.18 754.35
VPGTrans 47.003 19.113 26.333 64.323 69.543 46.179 76.224 30.474 76.764 38.276 42.626 44.816 74.149 46.003 56.044 2.509 760.30 714.86
MultiGPT 36.295 23.933 21.007 52.917 62.031 30.262 68.357 25.582 72.767 10.175 58.535 46.355 67.587 49.802 59.824 2.811 687.63 643.24
LaVIN-7B 75.113 25.673 22.892 39.322 40.752 26.847 71.415 25.735 69.977 29.461 37.221 33.061 60.087 50.407 60.487 12.345 680.73 632.17
LaVIN-13B 77.545 27.672 25.227 53.644 48.794 32.774 79.977 27.631 73.547 27.207 35.396 40.686 62.766 49.601 57.587 11.561 731.54 678.65
Lynx 38.287 22.725 19.225 64.327 67.719 42.796 74.777 22.372 73.727 10.282 46.006 47.0010 73.522 51.607 43.967 9.229 707.48 665.54
Fuyu-8B 45.715 27.0610 26.145 48.547 55.467 46.349 83.493 17.103 78.373 40.529 58.822 50.815 74.533 53.001 61.442 1.727 769.05 715.85
LaVIT 42.095 27.223 26.563 61.652 68.742 41.782 73.098 28.952 64.102 15.502 36.841 43.362 65.074 51.406 56.003 0.623 702.97 649.19
Qwen-VL-Chat 63.513 27.836 35.335 62.386 69.066 49.291 85.937 33.926 80.417 44.121 65.487 59.724 73.747 54.401 61.107 20.473 886.69 823.53
BLIP-2 74.171 28.005 35.2210 66.264 68.229 51.063 88.999 43.422 86.886 45.796 70.785 64.423 77.599 58.007 72.022 34.691 965.51 902.29
InstructBLIP 73.332 28.677 33.446 71.606 70.376 52.368 88.689 43.282 85.989 45.689 63.075 62.723 80.5310 58.209 73.107 36.722 967.73 905.62
Table 5: The mean relative gain, MRGM ($\uparrow$), for various models across instructions. Positive values indicate performance above the average, while negative values indicate performance below the average. The ‘Wins1’ and ‘Wins3’ columns display the number of tasks where a model achieves the highest aggregated performance and the top 3 aggregated performance, respectively. Text in ‘bold’, ‘underline’, and $\star$’ indicate the best scores, the sub-optimal scores, and the third-best scores, respectively. The notation “-” signifies that we do not assess LLMs for PuzzleVQA and MMMU datasets. The “Win1” and “Win3” results of LLMs are also not included.
Models ScienceQA PuzzleVQA MMMU MVSA-S MVSA-M TumEmo MOSI-2 MOSI-7 MOSEI-2 MOSEI-7 Twitter15 Twitter17 MASAD Hate Sarcasm MNRE Wins1 Wins3
ChatGPT 51.57 - - -2.81 -6.28 32.80 41.63 74.85 37.07 76.27 77.51 48.93 28.86 31.44 34.20 192.25 - -
LLaMA1-7B -50.97 - - -47.87 -49.57 -35.96 -22.66 -20.17 -25.81 -52.30 -37.98 -41.46 -46.55 -11.85 -22.24 -93.47 - -
LLaMA1-13B -51.84 - - -17.08 -11.25 -35.95 -19.93 -16.56 -12.02 -41.17 -22.56 -15.42 -19.51 -19.05 -32.17 -59.44 - -
LLaMA2-7B -37.93 - - -11.24 -8.86 -24.30 -31.27 -39.26 -9.84 -55.93 -38.14 -31.74 -36.70 -27.36 -18.75 -88.05 - -
LLaMA2-13B -47.79 - - -22.48 -21.48 -14.94 -20.75 -38.36 -26.87 -60.73 -35.68 -34.86 -44.32 -29.70 -28.18 -71.98 - -
Mixtral-AWQ 27.18 - - -3.26 -9.31 16.91 11.78 42.48 3.21 58.90 12.68 14.72 -4.69 20.96 28.54 142.48 - -
Gemma 13.00 - - 36.04 24.82 24.39 18.49 25.77 4.31 -12.56 10.11 19.31 9.50 21.96 10.43 120.24 - -
Flan-T5-XXL 63.30 - - 27.27 27.54 46.64 42.11 66.43 45.81 114.81 49.09 50.64 35.73 33.23 53.63 193.09 - -
Gemini-V 65.79 41.38 39.89 43.90 43.01 32.63 33.11 67.56 42.40 49.36 35.59 42.45 41.77 45.96 3.97 264.42 2 7
OpenFlamingo -58.69 -51.74 -47.50 -53.19 -48.84 -51.10 -35.87 -58.67 -40.51 -66.59 -52.68 -53.01 -51.20 -41.85 -47.47 -90.70 0 0
Fromage -69.73 -74.55 -85.98 -74.51 -85.75 -81.65 -70.07 -82.98 -74.15 -89.57 -83.14 -80.18 -75.18 -59.26 -54.39 -99.37 0 0
LLaVA-v0-7B -9.43 21.19 10.22 24.45 13.70 -43.60 -0.55 -0.51 3.98 -27.15 -24.06 -11.13 9.98 -55.75 -29.25 -79.68 0 0
LLaVA-v0-13B -3.41 33.74 28.05 37.15 31.96 -5.42 -0.48 -14.73 7.15 -8.21 -4.02 9.93 34.94 -33.90 -28.01 -76.11 0 0
LLaVA-v1.6-7B 36.14 13.17 27.05 14.98 30.02 22.74 28.39 6.49 34.04 -18.06 36.98 30.46 35.70 33.68 23.77 -15.09 0 0
LLaVA-v1.6-13B 48.30 20.65 16.24 23.36 19.47 46.72 35.57 38.27 22.98 72.55 46.66 37.21 36.85 40.21 31.31 55.97 0 1
MiniGPT4 13.93 43.05 23.04 46.26 26.33 45.01 17.41 24.29 26.56 16.10 7.62 19.08 31.61 24.48 8.88 -85.01 0 2
mPLUG-Owl -38.73 -17.61 -28.86 -16.13 -27.33 -30.05 -29.66 -30.29 -32.37 -49.21 -40.77 -39.54 -28.60 -3.56 -18.94 -74.00 0 0
mPLUG-Owl2.1 34.49 -5.04 14.97 -3.37 17.62 37.02 23.82 32.78 16.45 34.78 41.75 28.37 27.19 24.90 19.26 18.30 0 0
AdapterV2 20.91 42.27 41.71 45.14 45.48 2.61 -1.44 4.00 13.86 -24.76 -7.94 7.58 17.19 -9.36 -0.69 -71.76 0 4
VPGTrans -28.38 12.87 9.32 16.18 13.36 11.08 -15.41 -22.78 -9.53 -18.54 -27.94 -28.40 -10.38 -32.33 5.67 -90.78 0 0
MultiGPT -60.50 -52.95 -47.57 -52.77 -46.87 -51.39 -46.94 -49.45 -49.25 -73.79 -43.68 -46.95 -38.31 -28.71 -34.72 -91.09 0 0
LaVIN-7B -7.49 -43.67 -42.30 -42.17 -40.59 -54.40 -28.57 -38.72 -40.57 -51.82 -46.66 -50.47 -28.06 -6.67 7.94 -76.79 0 0
LaVIN-13B -2.81 -32.20 -45.98 -31.05 -44.38 -30.76 -8.13 -21.66 -13.81 -16.86 -40.28 -36.02 -16.56 -16.80 -23.97 -53.80 0 0
Lynx -51.03 -0.21 5.68 2.46 9.72 -21.40 -26.20 -62.03 -18.58 -72.15 -8.90 -4.54 4.67 -20.54 -39.09 -63.09 0 0
Fuyu-8B -21.47 -51.31 -33.77 -50.78 -32.50 18.30 14.21 -33.21 6.53 61.31 57.52 19.12 -19.56 22.49 15.60 -95.28 0 1
LaVIT -16.95 6.43 11.11 8.47 14.45 -7.85 -6.05 -5.16 -7.59 -43.01 -12.02 1.43 -2.22 6.60 -10.28 -98.07 0 0
Qwen-VL-Chat 32.27 17.06 26.06 20.02 29.53 41.36 13.89 22.51 3.97 78.03 61.58 45.57 22.21 20.21 30.63 45.97 0 1
BLIP-2 72.46 34.00 35.98 36.55 39.55 51.44 42.25 68.21 47.21 112.19 46.54 49.29 40.76 34.71 55.81 212.59 2 10
InstructBLIP 77.79 43.46 42.63 46.49 46.46 59.12 41.30 60.89 45.35 108.13 42.82 49.63 44.89 35.86 58.52 228.26 8 13
5 Experimental Results and Analysis
We evaluate the zero-shot performance of 31 LLMs (as detailed in Table 2) across 16 datasets (listed in Table 3), utilizing 10 distinct instructions per model (as shown in Figures 3, 4, and 5). The related results in a total of 4960 (31 $\times$ 16 $\times$ 10) experimental results, each yielding an accuracy score denoted as $acc_{mdi}$ from Eq. 1. Although most of the evaluated datasets could rely on text context for inference, PuzzleVQA and MMMU are exceptions, necessitating multimodal contexts. For ScienceQA, which primarily depend on the image modality to answer questions, we also explore the potential to extract knowledge from models using solely the image modality. Our comprehensive evaluation incorporates a range of metrics, including the Best Performance metric, the Mean Relative Gain metric, the Stability metric, and the Adaptability metric. Detailed analyses of these metrics and insightful conclusions related to them are presented in the subsequent sections.
5.1 Best Performance
We present the best performance ($A^{\tilde{i}}$), calculated using Eq. 2, achieved by different models across various multimodal datasets. The metric serves as benchmarks for the performance of different models on each dataset. Based on these results, we make the following observations:
5.1.1 Performance of Various Models on Full Datasets
We first present the best performance of various models across full datasets, as detailed in Table 4.
(1) Comparison between open-source models and closed-source models: While the Flan-T5 series, including Flan-T5, BLIP-2, and Instruct-BLIP, are strong open-source contenders, closed-source models such as Gemini-V typically outperform their open-source counterparts in multimodal reasoning tasks involving vision-text contexts. Notably, Gemini-V achieves the best performance across the most datasets; ChatGPT also delivers excellent performance using text-only contexts. However, it remains uncertain whether these closed-source models have been specifically trained on these datasets, a factor that could contribute to their superior performance. Another possible explanation for the observed discrepancies is that these closed-source models may be significantly larger than the open-source models evaluated. Unfortunately, we do not have specific information about the scale of the closed-source models. (2) Flan-T5 backbone vs. LLaMA backbone in the open-source models: Within open-source models, InstructBLIP stands out by securing top performance on three datasets and clinching the first position in the “Total” column, with BLIP-2 closely following in second place. Notably, both InstructBLIP and BLIP-2 utilize Flan-T5-XXL as their backbone model. Flan-T5-XXL is recognized as the best-performing large language model among the 8 LLMs evaluated (as shown in the first part of Table 4). It highlights the exceptional performance capabilities of Flan-T5-XXL. The superior performance of Flan-T5-XXL could be attributed to its Encoder-Decoder architecture, unlike the Decoder-only architecture of the LLaMA series models. The encoder module in Flan-T5-XXL offers significant advantages in multimodal representation learning, which is crucial for handling multimodal reasoning tasks with complex multimodal contexts. (3) LLMs vs. MLLMs: For the cumulative best performance of models across all datasets—excluding PuzzleVQA and MMMU, as indicated by the “Total” column—we observe that LLMs exhibit superior performance compared to certain MLLMs, such as Openflamingo, Fromage, LaVIT, and others. The phenomenon may be due to theses MLLMs tend to specialize in traditional visual language (VL) capabilities, including language generation, recognition, knowledge reasoning, spatial reasoning, and OCR, linked to the image modality. Meanwhile, these models might not sufficiently prioritize capabilities like multimodal alignment, multimodal fusion, and etc. Such capabilities are critical for effectively handling multimodal reasoning tasks that involve both vision and text contexts. (4) Different scale checkpoints: In the same model series, larger versions, evidenced by increased parameters such as in LLaMA-7B vs. LLaMA-13B, LLaVA-v0-7B vs. LLaVA-v0-13B, and LaVIN-7B vs. LaVIN-13B, generally outperform their smaller counterparts across most datasets. It suggests that models with more parameters and extensive training are likely to achieve superior performance. Furthermore, the latest model iterations, such as the mPLUG-OWL and LLaVA series, exhibit significant enhancements over previous versions, indicating that advancements in the novel architecture and more training substantially boost model efficacy. (5) Various instructions: On one hand, performance peaks vary within the same model when applied to different datasets, contingent on the instructions used. On the other hand, across various models, the optimal performance on a specific dataset often requires model-specific instructions. Despite its importance, the area of selecting the most suitable instructions for diverse datasets and models is under-researched and represents a promising avenue for future exploration. (6) More challenge tasks: Closed-source models like Gemini-V, significantly outperform open-source models on more challenging datasets, including MOSI-7 and MOSEI-7 which involve multiple classification categories, MNRE with long-tail effect, and more complex reasoning datasets such as PuzzleVQA and MMMU. There remains considerable potential for improvement in open-source models addressing these tasks. (7) With/Without text context on ScienceQA: ScienceQA primarily relies on vision context to answer questions. Accordingly, we conduct comparative experiments that exclude textual context, as detailed in the “Text context” part of Figure 3. The results of these experiments are shown in Figure 6. We observe that while most MLLMs demonstrate better performance on the ScienceQA dataset when textual context is included, the degree of improvement is limited. It suggests that the presence or absence of textual context has a limited impact on the performance of the ScienceQA dataset.
Figure 6: Comparison of experimental results in ScienceQA (with/without Text Context) for best performance, (measured by Accuracy), and Mean Relative Gain of MLLMs across all instructions. ‘w/’ means with and ‘w/o’ means without.
5.1.2 Full datasets vs. Subsets
Given the high cost and limited availability of GPT-4V and Claude3-opus, we draw inspiration from MultiPoint (Yang et al., 2023c) and GMP (Yang et al., 2023b) to tailor our approach. Specifically, we construct subsets by randomly sampling 10% of the data from the corresponding complete test datasets, ensuring that the distribution remains consistent with that of the original datasets. We then exclusively assess GPT-4V and Claude3-V on these subsets across all tasks, selecting corresponding experimental results from other models for comparison. The comparative analysis, focusing on the best performance metric between the full dataset and its subset, is depicted in Figure 7.
We employ the “LLMs-Total” metric to sum the best accuracy across all datasets, excluding PuzzleVQA and MMMU, to gauge overall performance for each LLM. Similarly, the “MLLMs-Total” metric sums the best performance across all datasets including PuzzleVQA and MMMU for assessing MLLM performance. The subplots “LLMs-Total” and “MLLMs-Total” indicate minimal performance differences between models on full datasets and their subsets, suggesting that our evaluation of models like GPT-4V and Claude3 on subsets is an accurate reflection of their overall performance across tasks. Moreover, GPT-4V consistently achieves superior performance on most tasks. It excels particularly in challenging datasets with extensive classification categories, such as MOSI-7 and MOSEI-7, datasets featuring long-tail effects like MNRE, and complex reasoning tasks including PuzzleVQA and MMMU, where it significantly outperforms other models. Notably, some open-source models, including InstructBLIP and BLIP2, also deliver competitive results on many subsets. An additional observation is that larger datasets tend to show less variation in experimental outcomes. The stability underscores the value of using subsets from large datasets, as they tend to provide more representative results.
Figure 7: Comparing results of the best performance metric between the full test dataset (Full dataset) and a subset of the test dataset (Subset). GPT-4V only has results on a subset, as indicated by the red dashed line (GPT-4V-Subset). The “Total” represents the sum of accuracy scores across all datasets for each model.
Table 6: The mean relative gain, MRGI ($\uparrow$), for different instructions (# 1, …, # 10) across models.
Instructions ScienceQA PuzzleVQA MMMU MVSA-S MVSA-M TumEmo MOSI-2 MOSI-7 MOSEI-2 MOSEI-7 Twitter15 Twitter17 MASAD Hate Sarcasm MNRE Wins1 Wins3
LLMs
# 1 -31.20 - - 47.67 40.38 18.59 23.92 32.41 25.20 34.71 11.20 27.72 32.72 19.42 31.62 29.26 2 11
# 2 9.74 - - 42.31 33.96 23.55 25.76 34.50 29.50 42.35 15.21 27.62 44.77 22.22 19.56 -15.44 6 11
# 3 -1.13 - - -36.65 -21.94 -31.35 6.66 -1.67 -11.22 -19.67 53.96 22.71 -10.58 3.66 -18.07 25.90 1 1
# 4 -24.65 - - -40.21 -35.02 -9.68 -32.93 -17.90 -28.79 -16.87 -53.47 -56.06 -56.65 2.49 33.41 -51.73 1 1
# 5 39.31 - - 1.05 14.52 10.84 20.52 16.44 -1.31 7.29 52.74 35.01 13.41 23.91 0.79 112.35 3 7
# 6 28.30 - - 5.69 6.76 20.59 -12.17 -3.57 -6.68 9.22 3.85 5.85 14.03 -36.43 -12.51 -59.13 0 3
# 7 43.89 - - 23.58 23.30 20.59 3.27 -31.00 17.04 -25.70 18.35 27.72 43.17 22.49 12.50 70.00 1 10
# 8 -36.80 - - 16.51 16.74 2.93 -5.61 7.73 11.31 4.18 -43.88 -39.03 -37.97 1.89 -13.93 -53.31 0 0
# 9 2.95 - - -29.78 -38.06 -16.58 9.04 10.31 -3.63 8.71 -15.77 -8.44 -0.69 -41.83 -23.02 -30.17 0 0
# 10 -30.41 - - -30.18 -40.64 -39.48 -38.45 -47.25 -31.44 -44.21 -42.18 -43.11 -42.21 -17.82 -30.36 -27.71 0 0
MLLMs
# 1 -26.34 -20.00 -15.50 14.44 8.51 23.89 22.92 26.50 11.87 25.10 2.70 10.17 1.88 4.46 11.49 3.27 1 7
# 2 11.39 10.41 9.00 24.41 27.99 23.30 21.64 44.07 19.73 37.61 16.01 21.46 17.31 20.38 14.34 0.44 6 12
# 3 16.41 21.68 11.02 -16.28 -10.04 -13.92 -3.78 -16.97 -2.16 -12.36 5.19 -2.09 -3.99 -7.33 -11.89 27.15 1 4
# 4 -17.63 -23.99 -21.69 -7.68 -5.64 -3.19 -9.41 -4.00 -8.62 -1.11 -11.44 -11.57 -12.21 -1.62 5.36 -26.03 0 0
# 5 37.68 22.68 25.35 -16.39 -12.94 -20.90 6.80 -4.75 4.26 -12.65 18.41 8.96 3.17 6.88 1.35 17.54 4 6
# 6 -7.35 -8.49 -1.33 10.87 15.79 3.07 -13.52 -1.87 -0.87 10.96 14.29 12.63 11.64 -1.18 3.64 5.17 0 6
# 7 29.09 14.62 18.23 23.92 17.63 6.24 26.35 -14.03 30.65 -6.35 11.31 11.23 25.15 11.00 14.44 2.80 4 12
# 8 -26.78 -11.31 -13.13 -14.57 -18.06 -6.63 -28.67 -16.88 -31.77 -27.00 -30.44 -28.99 -27.11 -17.22 -18.86 -17.57 0 0
# 9 6.32 10.29 3.45 -0.62 -1.88 0.49 0.22 9.74 1.60 8.66 -5.63 -3.45 -0.97 -7.41 -10.42 -6.23 0 1
# 10 -22.78 -15.89 -15.39 -18.11 -21.36 -12.35 -22.54 -21.82 -24.69 -22.86 -20.40 -18.35 -14.87 -7.95 -9.45 -6.54 0 0
Table 7: The stability, $S^{\mathcal{M^{{}^{\prime}}}}$ ($\downarrow$), of various models with excellent performance across instructions.
Models ScienceQA PuzzleVQA MMMU MVSA-S MVSA-M TumEmo MOSI-2 MOSI-7 MOSEI-2 MOSEI-7 Twitter15 Twitter17 MASAD Hate Sarcasm MNRE Wins1
ChatGPT 4.36 - - 13.44 8.37 3.61 1.56 2.17 5.90 3.57 2.53 3.20 6.32 6.9 9.75 6.92 -
Flan-T5-XXL 0.57 - - 10.84 8.99 4.62 0.73 3.73 0.38 2.97 9.63 3.09 0.52 0.93 1.75 2.19 -
Gemini-V 15.69 0.39 4.97 11.35 8.71 6.59 8.41 7.40 3.91 6.46 7.83 5.53 7.38 10.24 4.93 5.81 1
BLIP-2 5.73 0.99 1.29 5.46 7.86 3.05 0.31 3.62 0.42 3.43 9.04 3.55 0.53 0.69 1.52 1.89 2
InstructBLIP 0.73 1.46 1.07 4.15 4.44 0.62 0.62 5.02 0.51 4.24 7.02 3.29 0.92 0.65 1.27 3.90 6
Table 8: The stability, $S^{\mathcal{I^{{}^{\prime}}}}$ ($\downarrow$), of different instructions with excellent performance across models.
Instructions ScienceQA PuzzleVQA MMMU MVSA-S MVSA-M TumEmo MOSI-2 MOSI-7 MOSEI-2 MOSEI-7 Twitter15 Twitter17 MASAD Hate Sarcasm MNRE Wins1
LLMs
# 1 24.65 - - 9.41 10.21 8.21 9.14 7.77 9.06 11.77 12.28 8.48 9.65 9.68 7.34 14.23 3
# 2 17.78 - - 8.61 12.90 6.00 9.81 6.75 8.52 10.93 11.91 8.60 5.33 8.62 11.92 16.79 7
# 3 24.70 - - 26.37 19.58 19.68 15.66 11.29 19.80 12.83 11.79 11.82 24.92 16.93 17.92 12.15 0
# 5 15.76 - - 16.93 12.43 11.59 11.84 8.00 19.33 11.25 13.05 7.55 16.82 6.43 11.65 11.90 1
# 6 14.37 - - 13.88 16.81 9.14 24.00 11.99 22.62 14.26 12.77 12.84 13.54 21.96 20.72 13.78 1
# 7 15.65 - - 10.01 13.41 7.11 17.42 17.37 10.07 15.85 11.67 9.43 5.36 8.71 16.96 11.48 2
MLLMs
# 1 22.68 7.83 10.48 13.79 16.72 11.40 11.28 10.36 16.05 13.39 15.22 15.70 21.11 15.06 14.76 12.07 3
# 2 16.11 6.29 8.32 15.35 16.86 11.44 10.81 8.56 14.89 12.04 12.21 12.22 14.94 10.75 12.40 12.00 6
# 3 19.72 5.64 9.20 24.22 23.60 16.58 21.93 10.47 20.92 11.24 15.76 16.07 23.10 17.60 20.48 10.10 2
# 5 16.45 6.50 7.74 22.66 21.04 16.93 19.20 10.92 17.88 10.75 18.29 16.58 20.90 15.65 19.49 11.70 2
# 6 17.67 7.97 9.27 17.20 17.70 13.74 26.78 12.10 23.78 14.47 15.03 14.88 19.08 15.64 13.31 11.40 0
# 7 14.02 6.45 8.85 17.14 19.98 12.49 12.19 10.83 12.64 13.95 13.88 12.67 13.03 13.51 13.89 11.67 3
Figure 8: The Top-K Hit Ratio, $GHR@K$ ($\uparrow$), for various instructions (# 1, …, # 10) across different models on all datasets. The “Total” represents the sum of Global Top-K Hit Ratio scores across all datasets for each model.
5.2 Mean Relative Gain
While the best performance metric offers insight into the upper limit of capabilities for each model, a comprehensive assessment on the each full dataset requires a broader perspective.
5.2.1 Mean Relative Gain of Models
To assess the comprehensive performance of each model, we calculate the mean relative gain for each model across all instructions on each dataset, as detailed in Eq. 3. The results are tabulated in Table 5. The metrics “Wins1” and “Wins3” denote the number of tasks where a model or an instruction achieves the highest and top-three aggregated performances, respectively. Similar to the best model performance metric, models equipped with the Flan-T5-XXL backbone, such as InstructBLIP, BLIP-2, and Flan-T5-XXL itself, exhibit strong performance. Closed-source models also show excellent results.
5.2.2 Mean Relative Gain of Instructions
We continue to evaluate the overall efficacy of each instruction by calculating the mean relative gain (MRGI) across all models, as detailed in Eq. 4. The outcomes are presented in Table 6. Due to the different evaluation datasets for LLMs and MLLMs—with LLMs excluding PuzzleVQA and MMMU and MLLMs including these datasets—we compute MRGI separately for each group. Our analysis reveals that Instructions # 1, # 2, # 5, and # 7 generally perform better, with their detailed designs illustrated in Figures 3 and 4. Notably, Instructions # 5 and # 7, which evolve from simple “Question-Answer” format of # 2, show superior performance. The effectiveness likely stems from the prior training of models on related QA tasks, where specific terms like “Question” and “Answer” enhance model responsiveness. Furthermore, the inclusion of option terms in Instructions # 5 and # 7 appears to positively influence performance on specific datasets.
5.3 Stability
Stability is a key metric for assessing both models and instructions on each full dataset. Notably, some models display consistent performance across all instructions, resulting in high stability, albeit at lower performance levels. Consequently, our evaluations prioritize the stability of models and instructions that show consistently positive performance.
5.3.1 Stability of Models
We exclusively use Eq. 5 to calculate stability metrics for models demonstrating strong overall performance, such as ChatGPT, Flan-T5-XXL, Gemini-V, BLIP2, and InstructBLIP, with the results displayed in Table 7. In terms of stability among open-source models, the Flan-T5 series exhibits minimal fluctuations across different instructions, indicating a lower sensitivity to instruction variability. Compared to BLIP2, InstructBLIP enhances stability by incorporating instruction tuning, effectively reducing the sensitivity of models to varying instructions in multimodal reasoning tasks with vision-text contexts. The improvement suggests that instruction tuning is a valuable strategy for enhancing model stability.
5.3.2 Stability of Instructions
We assess the stability of instructions that achieve a top-three average relative gain on at least one dataset from LLMs and MLLMs using Eq. 6, as detailed in Table 8. Notably, Instruction # 2 exhibits the highest stability and also demonstrates strong aggregated performance. The superior stability likely stems from the prevalent training of models on the “Question-Answer” data format. The finding provides valuable insights for the design of future instructions.
5.4 Adaptability
Models are trained with distinct pre-training settings, leading to varying preferences towards different instructions. We quantify the adaptability using the Global Top-K Hit Ratio, as detailed in Eq. 7. To thoroughly assess the adaptability of various models to different instructions across all full datasets, we have categorized all models into nine groups based on their language backbones. These groups include Closed-source, LLaMA LLMs, Other LLMs, MLLMs based on LLaMA, MLLMs based on Vicuna, MLLMs based on Flan-T5, LLaVA Series, LaVIT Series, and MLLMs based on other language backbones. The comprehensive results, covering all datasets, are illustrated in Figure 8. It’s important to note that evaluations for LLMs exclude the AlgoPuzzleVQA and MMMU datasets, while MLLMs include the two datasets.
5.4.1 Total Adaptability
Overall, instructions # 2 and # 7 demonstrated the highest adaptability among all models as the “Total” subplot shown in Figure 8, indicating that these instructions should be prioritized in the design of future models.
5.4.2 Adaptability for Different Models
We advert that different models tend to perform better with specific instructions. (1) Closed-source models: Closed-source models exhibit their poorest performance on Instructions # 6 and # 9, suggesting that specific trigger words such as “AI:”, “Human:”, and “### Input” significantly impair their performance. (2) Same models with different version: The LLaMA model series, which includes LLaMA1-7B, LLaMA1-13B, LLaMA2-7B, and LLaMA2-13B, consistently shows strong performance on Instructions # 1 and # 2, indicating similar preferences across different models. Similarly, the LaVIN model branch excels with Instructions # 2 and # 7. In contrast, the LLaVA model branches demonstrate more varied preferences, as models of different sizes optimize performance using different instructions. The variation in the LLaVA models may be attributed to differences in training methodologies across different scales. (3) Various backbones of LLMs: We further categorize additional models into four groups based on large language model backbones. Related groups are listed that models using LLaMA, Vicuna, Flan-T5, and others as their LLMs. We note that models based on different backbones exhibit varied performance across diverse instructions. For instance, most models utilizing the LLaMA backbone excel on Instructions # 1, # 2, and # 7. However, certain models, such as OpenFlamingo and mPLUG-OWL, struggle with specific instructions, achieving an adaptability score of zero on # 4, among others. Models with Vicuna backbones generally underperform on Instructions # 8 and # 1. Conversely, the Flan-T5 series, which includes Flan-T5, BLIP-2, and InstructBLIP, shows consistent adaptability trends, performing well on Instructions # 9 and # 10 but less so on Instructions # 1 and # 8. Models leveraging other LLM backbones display optimal performance on a range of instructions, with Fromage, in particular, experiencing significant performance variability across different instructions.
Our observation reaffirms that the training of each model involves carefully designed instructions, including specific trigger words. Consequently, to achieve the best performance, it is essential to tailor instructions for each model based on the prompts used during its training. Furthermore, MM-InstructEval offers crucial guidance and direction for crafting instructions that maximize the performance of both existing and emerging models.
6 Conclusion
We conduct comprehensive evaluations and establish benchmarks for 31 models, including 23 MLLMs, using 10 instructions across 16 diverse multimodal reasoning datasets with vision-text contexts. Our framework, MM-InstructEval, introduces a range of multi-perspective metrics to assess these tasks comprehensively. These metrics include the Best Performance metric, the Mean Relative Gain metric, the Stability metric, and the Adaptability metric. Through the comprehensive evaluation, we have drawn several important conclusions: (1) Models utilizing the Flan-T5-XXL backbone, which is based on an Encoder-Decoder architecture, consistently outperform those from the LLaMA series, which employ a Decoder-only architecture, in our multimodal reasoning evaluations. (2) Instructions formatted in a “Question-Answer” style tend to yield better performance. (3) There is a notable variation in model preferences for specific instructions across different models. (4) Excellent closed-source models, such as GPT-4V and Claude3, demonstrate superior overall performance across all tasks, excelling in challenging datasets. The study about Multimodal Large Language Models is a rapidly evolving field. Our research lays the groundwork for new exploratory directions and assesses the zero-shot performance of MLLMs in complex multimodal reasoning tasks. Looking forward, we aim to explore the performance of MLLMs in in-context learning scenarios in future studies.